OpenStudy (anonymous):

what is the integral of cos(x)*e^x? plz help was just testing the site before

OpenStudy (unklerhaukus):

integration by parts, recursive

OpenStudy (anonymous):

?? tried doing by parts but it ddidn't seem to work

OpenStudy (anonymous):

by parts would definitley work

OpenStudy (unklerhaukus):

do by parts twice, and then notice that your left with the original thing left to integrate

OpenStudy (unklerhaukus):

Show some working

OpenStudy (anonymous):

it is very messy... are there any other ways I can try instead?

OpenStudy (anonymous):

you could do some trig which would be abstract and would make it worse

OpenStudy (anonymous):

I'd rather take that

OpenStudy (anonymous):

you would still have to integrate by parts

OpenStudy (anonymous):

just set it up

OpenStudy (anonymous):

u = e^x du = e^x dv = cos(x) v= ?